### 23 Nov 2023 v1

# HACD: Hand-Aware Conditional Diffusion for Monocular Hand-Held Object Reconstruction

Abstract
Reconstructing hand-held objects from a single RGB image without known 3D object templates, category prior, or depth information is a vital yet challenging problem in computer vision. In contrast to prior works that utilize deterministic modeling paradigms, which make it hard to account for the uncertainties introduced by hand- and self-occlusion, we employ a probabilistic point cloud denoising diffusion model to tackle the above challenge. In this work, we present Hand-Aware Conditional Diffusion for monocular hand-held object reconstruction (HACD), modeling the hand-object interaction in two aspects. First, we introduce hand-aware conditioning to model hand-object interaction from both semantic and geometric perspectives. Specifically, a unified hand-object semantic embedding compensates for the 2D local feature deficiency induced by hand occlusion, and a hand articulation embedding further encodes the relationship between object vertices and hand joints. Second, we propose a hand-constrained centroid fixing scheme, which utilizes hand vertices priors to restrict the centroid deviation of partially denoised point cloud during diffusion and reverse process. Removing the centroid bias interference allows the diffusion models to focus on the reconstruction of shape, thus enhancing the stability and precision of local feature projection. Experiments on the synthetic ObMan dataset and two real-world datasets, HO3D and MOW, demonstrate our approach surpasses all existing methods by a large margin.
1 Introduction
Figure 1: We present hand-aware conditioning, utilizing both semantic and geometric embeddings to model hand-object interaction. We further propose hand-constrained centroid-fixed diffusion, where the centroid of the partially denoised point cloud at each step is constrained by the hand vertices and kept fixed throughout the whole diffusion and the reverse process.
Modelling hand-object interactions has a wide range of applications across various domains, including AR/VR [38, 54, 55] and robotics [75, 24, 72]. Significant progress has been made in this task [38, 72, 4], in which single-view hand-held object reconstruction draws particularly great attention [30, 37, 73]. Reconstructing hand-held objects from a single RGB image is a highly challenging and ill-posed problem. Suffering from the lack of real-world data and the ambiguity caused by hand- and self-occlusion, the performance of single-view hand-held object reconstruction remains limited.
Most existing works rely on Signed Distance Fields (SDFs) [37, 73] to represent the object shapes. Though effective under favorable conditions, such technique tends to result in over-smoothed and undetailed reconstructions [51, 10]. Moreover, prior works [30, 37, 73] typically utilize a deterministic modeling paradigm, which makes it hard to reason about the uncertainties introduced by hand- and self-occlusion. Recently, probabilistic point cloud denoising diffusion models [43, 80, 44] have shown to be effective for the task of single-view object reconstruction. First, compared to works employing surface-based representations like SDFs, approaches reconstructing point clouds enjoy better capabilities in overcoming artifacts such as noise, sparsity, and irregularity [10]. Second, the probabilistic nature of denoising diffusion models is particularly beneficial for modeling uncertainty and indeterminacy. Therefore, in this work, we propose to leverage these probabilistic point cloud denoising diffusion models to conduct hand-held object reconstruction from a single RGB image.
Directly employing diffusion models in single-view hand-held object reconstruction, however, faces two main problems. First, diffusion models for 3D reconstruction typically assume isolated and unoccluded objects in the images [80, 44, 43] with self-occlusion properly modeled, while in hand-held object reconstruction tasks, hand occlusion needs to be modeled as well. In fact, the shape of the hand-held object is constrained by and highly related to the hand semantically and geometrically. Second, the object centroid is generally constrained by the hand vertices during the diffusion and the reverse processes. In existing diffusion models, the centroid of a partially denoised point cloud is not controlled and left variable, thus the centroid can easily deviate and move, for example, to the back side of the hand or even intersect with the hand, leading to physically implausible results. Additionally, the centroid bias causes the misalignment of the semantic features [19] and can thus have adverse effects on the object reconstruction quality.
To solve the aforementioned problems, we present HACD, a Hand-Aware Conditional point cloud denoising Diffusion probabilistic model for single-view hand-held object reconstruction (Fig. 1). To fully utilize the hand prior, we introduce hand-aware conditioning to model hand-object interaction from both semantic and geometric perspectives. Via extraction of local features from object and hand, we acquire a unified hand-object semantic embedding, which compensates for the semantic feature deficiency induced by the hand-occluded object region. Further, at each step, we explicitly encode the relative position between hand articulation and the partially denoised point cloud to obtain a geometric hand articulation embedding. To fuse the aforementioned semantic and geometric embeddings, we propose a novel dual-stream conditioning to separately predict noise features in two branches and then estimate the final noise from the integrated feature. Additionally, a hand-constrained centroid fixing scheme is proposed to ensure that the centroid of the partially denoised point cloud does not diverge during the diffusion as well as the reverse process. In particular, we use a small and efficient neural network to estimate the hand-constrained object centroid to properly guide the reverse diffusion. As a consequence, the model only needs to take care of the simpler shape diffusion, rather than having to simultaneously account for shape and centroid. Moreover, leveraging the hand vertices constraint together with our centroid fixing can further enhance the precision and stability of our semantic and geometric features. Our contributions can be summarized as follows.
  • We present HACD, a hand-aware conditional point cloud diffusion model for monocular hand-held object reconstruction, demonstrating superior performance in modeling the uncertainties induced by hand- and self-occlusion.
  • We introduce a unified semantic hand-object embedding and a geometric hand articulation embedding to model semantic and geometric aspects of the shown hand-object interaction, which is subsequently utilized within our proposed dual-stream conditioning.
  • We further propose a hand-constrained centroid fixing scheme, utilizing the hand vertices prior to prevent the centroid of the partially denoised point cloud from diverging during diffusion and its reverse processes.
Experiments on the synthetic ObMan [30] dataset and two real-world datasets, namely HO3D [29] and MOW [4], demonstrate that HACD can surpass all existing methods by a large margin.
2 Related Work
Hand Pose estimation. Hand pose estimation methods from RGB(-D) images can be primarily categorized into model-free and model-based methods. Model-free approaches typically detect 2D keypoints, which are then lifted to 3D joint positions [57, 58, 82, 47, 50, 34, 48], while model-based approaches exploit statistical models like MANO [59] with low-dimensional parameter spaces [63, 1, 79, 81, 60]. Compared with model-free methods, the model-based methods exhibit better robustness to occlusions and domain discrepancies. Therefore, our method utilizes an off-the-shelf model-based approach [60] to predict hand poses, which are then further utilized for shape inference of hand-held objects.
Single-view Object Reconstruction. Object reconstruction is a long-standing problem in the computer vision community. Traditionally, varieties of works focus on reconstruction from multiple-view geometric techniques [61, 76, 15, 16]. Recently, advanced by the powerful representation capacity of deep learning, single-view-based methods for object reconstruction have shown promising results [44, 11, 70, 71], in spite of the highly ill-posed setting. Early learning-based methods propose to learn category-specific networks [36, 35, 67], while more recent works try to learn a generalizable model across multiple categories for either meshes [28, 26], voxels [25, 11, 69], point clouds [22, 39], or implicit representations such as NeRFs [45] and SDFs [51, 7]. In this work, we focus on the difficult yet very important problem of learning hand-held object reconstruction from a single RGB image [30], putting a particular emphasis on modeling the impact of the hand occlusion.
Hand-Held Object Reconstruction. Reconstructing hand-held objects is very challenging due to intricate mutual occlusions and self-occlusions. Earlier works opt for the simplified task of 6DoF object pose estimation [17, 64, 66, 29, 31], assuming known object templates. To better utilize hand-object interactions, some studies reason about hand and object poses by means of implicit feature fusion [6, 40, 62, 66], exploiting explicit geometric constraints [12, 2, 27, 4, 18, 77, 78], or enforcing of physical constraints [68, 52]. Increasing attention has recently been paid to model-free hand-held object reconstruction, as it is a more applicable setting. Exemplarily, while [30] explicitly predicts the object mesh, [37, 8, 9] leverage implicit 3D representations such as signed distance functions for reconstructing the shape of the hand and object. However, all these methods follow a deterministic approach, oftentimes resulting in low-quality reconstruction results for occluded and invisible parts, especially in the ill-posed single-view setting. We instead draw inspiration from recent advances in probabilistic 3D generation [33, 80, 19, 44] and additionally leverage hand-articulation priors to achieve higher-fidelity hand-held object reconstruction.
[…]
3 Method
Figure 2: Architecture of HACD. (I) Given a single-view RGB image, we first predict the hand pose $\phi_{H}$ and camera view $\phi_{C}$ by an off-the-shelf network. (II) The object centroid $\widehat{\mathcal{M}}$ is then estimated by our simple yet efficient hand-constrained centroid prediction network. (III) We further introduce a centroid-fixed diffusion network, which always keeps the centroid of partially denoised point cloud fixed at the predicted centroid $\widehat{\mathcal{M}}$ during the reverse process. (IV) Both the unified semantic hand-object embedding and geometric hand articulation embedding are utilized as condition in the diffusion and the reverse process, which is then fused by a dual-stream conditioning scheme.
Given a single-view RGB image $I$ capturing a hand-held object, we aim at reconstructing the 3D point cloud of the object. As shown in Fig. 2, we first utilize an off-the-shelf method [30, 60] to predict hand parameters $\phi_{H}$ and camera view $\phi_{C}$, where $\phi_{H}$ is defined using the MANO [59] hand model with 45DoF joint parameters and $\phi_{C}$ represents the 6DoF pose of the hand wrist in the world reference system. A hand-constrained centroid prediction network is then employed to predict the object centroid $\widehat{\mathcal{M}}$ given the estimated hand vertices. Eventually, we use our centroid-fixed conditional point cloud diffusion model to reconstruct the object point cloud. Compared with existing diffusion models, our model can thus fully concentrate on the reconstruction of the shape, leading to an improved performance. We further leverage semantic as well as geometric hand priors during diffusion. The unified semantic hand-object embedding and geometric hand articulation embedding are further used within our dual-stream conditioning. Due to the probabilistic nature of diffusion models together with our explicit modeling of hand-object interaction, HACD shows superior performance in handling uncertainties arising from hand- and self-occlusion.
In the following sections, we first introduce the fundamentals of conditional point cloud denoising diffusion models (Sec. 3.1). We then explain our novel hand-aware conditioning (Sec. 3.2), before describing our hand-constrained centroid-fixing scheme (Sec. 3.3).
3.1 Conditional Point Cloud Denoising Diffusion
We formulate single-view hand-held object reconstruction as conditional point cloud denoising diffusion, consisting of two Markov chains called the diffusion process and the reverse diffusion process. Suppose that we have a target point cloud with $N$ points $X_{0}\in\mathbb{R}^{3N}$ from the conditional distribution $q(X|z)$, where $z=I_{\phi_{C},\phi_{H}}$ is the input RGB image capturing the hand-held object with the corresponding camera view $\phi_{C}$ and the hand pose $\phi_{H}$. For the diffusion process, we gradually add Gaussian noise to the target point cloud at different levels $t\in\{1,\cdots,T\}$ as
$q(X_{t}|X_{t-1},z)=\mathcal{N}(X_{t}|z;\sqrt{1-\beta_{t}}X_{t-1}|z,\beta_{t}\mathbf{I}).$ (1)
Notice that with a fixed variance schedule $\{\beta_{t}\}_{t=0}^{T}$, $X_{t}$ can be simply expressed by $X_{0}$ according to
$q(X_{t}|X_{0},z)=\mathcal{N}(X_{t}|z;\sqrt{\bar{\alpha}_{t}}X_{0}|z,(1-\bar{\alpha}_{t})\mathbf{I}),$ (2)
where $\alpha_{t}=1-\beta_{t},\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$. Therefore, we can reparameterize $X_{t}$ as a linear combination of $X_{0}$ and a noise variable $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ as follows
$X_{t}=\sqrt{\bar{\alpha}_{t}}X_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon.$ (3)
Starting with a point cloud sample $X_{T}$ from random Gaussian noise, the reverse process iteratively samples from $q(X_{t-1}|X_{t},z)$ to remove the added noise from the diffusion process. To approximate this reverse process, we train a conditional denoiser $\mathcal{D}_{\theta}(X_{t},t,z)$ to learn the distribution $q(X_{t-1}|X_{t},z)$ by
$p_{\theta}(X_{t-1}|X_{t},z)=\mathcal{N}(X_{t-1};\mu_{\theta}(X_{t},t,z),\sigma_{t}^{2}\mathbf{I}),$ (4)
$\mu_{\theta}(X_{t},t,z)=\frac{1}{\sqrt{\alpha_{t}}}(X_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(X_{t},t,z)),$
where $\mu_{\theta}$ is the estimated mean.
3.2 Hand-Aware Conditioning
The essence of the conditional point cloud denoiser $\mathcal{D}_{\theta}(X_{t},t,z)$ is the modeling of the condition $z$ so to fully utilize the information of the input image along with the corresponding camera view and hand pose. Our key insight is that the hand can offer both semantic and geometric priors to facilitate object reconstruction. Therefore, a unified semantic hand-object embedding and a geometric hand articulation embedding are proposed in an effort to well model the hand-object interaction.
Unified Semantic Hand-Object Embedding. The 2D image feature is a vital component for object reconstruction. Unlike prior works that extract a global embedding for the image [11, 70, 71], we choose to compute unique semantic features for each point. Specifically, we first extract image features $\mathcal{F}\in\mathbb{R}^{H\times W\times C}$ by a standard 2D network such as ResNet [32] or ViT [21], with $C$ being the number of feature channels. Instead of sharing the same global features among all points, we project the extracted image features onto the partially denoised point cloud at each diffusion step. Therefore, we design a corresponding projection function $\pi:\mathbb{R}^{H\times W\times C}\rightarrow\mathbb{R}^{N\times C}$, where $N$ denotes the number of points in the point cloud. We implement $\pi$ using the efficient point cloud rasterizer [56]. For every individual point within the point cloud, we obtain semantic features corresponding to the pixel onto which the point is rasterized. Subsequently, the pointwise semantic features at timestep $t$ can be represented as $X_{t}^{O}=\pi(\mathcal{F},\phi_{C},X_{t})\in\mathbb{R}^{N\times C}$.
However, due to the inevitable occlusion induced by the hand, it is very challenging to reconstruct the occluded part of the object as only a single view is provided. Due to the incomplete object semantic features, relying solely on the local features of object points alone is insufficient to conduct precise hand-held object reconstruction. On the other side, the hand information should not be fully ignored as the hand naturally contains information on the shape of the hand-occluded object region. Therefore, we apply $\pi$ to both object and hand points to obtain $X_{t}^{HO}=\pi(\mathcal{F},\phi_{C},[X_{t},P_{H}])\in\mathbb{R}^{(N+N_{H})\times C}$, which serves as a unified semantic hand-object embedding. Thereby, $N_{H}$ denotes the number of the hand mesh vertices $P_{H}$, which can be obtained from the estimated hand pose $\phi_{H}$ together with the MANO hand model [59]. Noteworthy, we also apply an extra one-hot encoding to indicate whether the points belong to the object or the hand, hence making $X_{t}^{HO}\in\mathbb{R}^{(N+N_{H})\times(C+1)}$.
Geometric Hand Articulation Embedding. Furthermore, the object shape is also highly constrained by the hand and very geometrically related to the hand articulation. In the core, the positions and directions of the hand joints in contact with the object, are supposed to be in accordance with the object surface for a stable hold. Therefore, knowing these properties can serve as a strong prior to enhance reconstruction quality. In practice, we explicitly encode hand-object interaction by transforming the partially denoised point cloud at every step to each hand joint frame.
Given the partially denoised point cloud $X_{t}$ and hand parameter $\phi_{H}$, we first compute the rotation $R_{j}$ and translation $T_{j}$ of each joint $j$ with respective to the hand wrist using forward kinematics given the hand model. We then transform $X_{t}$ from the hand wrist coordinate system to each hand joint frame via $X_{t}^{j}=R_{j}X_{t}+T_{j}\in\mathbb{R}^{N\times 3}$ to encode the hand articulation onto object points. Finally, the hand articulation embedding $X_{t}^{A}\in\mathbb{R}^{N\times J}$ is calculated via concatenation and flattening of $X_{t}^{j}$, with $J$ equaling 45 in our experiments as 15 hand joints are utilized.
Dual-Stream Conditioning. One possible approach is to directly concatenate the unified semantic hand-object embedding $X_{t}^{HO}$ and geometric hand articulation embedding $X_{t}^{A}$ as $z=[X_{t}^{HO},X_{t}^{A}]$ and use it as condition. However, forcibly integrating semantic and geometric embeddings may lead to mutual interference. Instead, we propose a dual-stream conditioning strategy to fuse $X_{t}^{HO}$ and $X_{t}^{A}$. In essence, we process $z_{1}=X_{t}^{HO}$ and $z_{2}=X_{t}^{A}$ separately through two distinct network branches, resulting in two features of identical dimensions, as shown in Fig. 2. The noise $\epsilon_{\theta}$ is then predicted from the concatenated feature.
3.3 Hand-Constrained Centroid Fixing
Existing works commonly obtain the camera view $\phi_{C}$ and the hand pose $\phi_{H}$ using standard hand pose estimation methods [30, 60]. Without knowing the object pose during inference, we would need to denoise the points in the hand wrist coordinate system. However, directly using a diffusion model to simultaneously learn the object’s shape and pose is very difficult [19]. Thus, we propose a hand-constrained centroid fixing scheme to ease learning.
Centroid-Fixed Diffusion. We first reduce the problem of object pose estimation to centroid prediction, by defining a new object coordinate system. Essentially, while the origin of the object coordinate system is located at the centroid of the object, the axis orientation is shared with the hand wrist frame. Notice that during training, we can directly use the ground-truth object centroid $\mathcal{M}$ to constrain the object point cloud. Therefore, we harness $\mathcal{M}$ to stabilize the diffusion process and make the point cloud fixed at $\mathcal{M}$. In particular, during centroid-fixed diffusion, we re-center the object point cloud to $\mathcal{M}$ and guarantee the added noise to be of zero-mean by
$X_{0}\sim q(X_{0}),$ $~{}X_{0}\leftarrow X_{0}-\bar{X_{0}}+\mathcal{M},$ (5)
$\epsilon\sim\mathcal{N}(\mathbf{0},$ $~{}\mathbf{I}),\epsilon\leftarrow\epsilon-\bar{\epsilon}.$
Noteworthy, by keeping the object centroid unchanged, the misalignment error in semantic feature projection due to centroid movements could also be alleviated. Thus, the training behavior becomes more stable as well.
Centroid Prediction for Reverse Diffusion. During the reverse diffusion process, we propose a simple yet effective hand-constrained network $\mathcal{G}$ to predict the translation of the object w.r.t. the hand wrist coordinate system. Given the input RGB image $I$ along with the corresponding hand pose $\phi_{H}$ and camera view $\phi_{C}$, the predicted object centroid is obtained by
^$\widehat{\mathcal{M}}=\mathcal{G}(I_{\phi_{C},\phi_{H}}).$ (6)
As shown in Fig. 2 (II), $\mathcal{G}$ first encodes the hand vertices $P_{H}$ using PointNet [53] to constrain the object centroid in 3D space. Subsequently, the global hand features are combined with the image features extracted from a pre-trained ResNet-18 [32], followed by two parallel multilayer perceptrons (MLPs) to respectively output the 3D and 2D object centroids.
[…]
$\epsilon_{\theta}\leftarrow\epsilon_{\theta}-\bar{\epsilon}_{\theta},$ (7)
[…]
After the denoising process, we can directly obtain the reconstructed object point cloud by taking $X_{0}$.
3.4 Training Objectives
We adopt Point-Voxel CNN (PVCNN) [41] for our diffusion model $\mathcal{D}_{\theta}$, as it has demonstrated great results in 3D object reconstruction tasks [80, 19, 44]. The point cloud along with its corresponding condition $[X_{t},X_{t}^{Z}]\in\mathbb{R}^{(3+C^{\prime})N}$ is thereby fed into $\mathcal{D}_{\theta}$ to predict the noise $\epsilon_{\theta}\in\mathbb{R}^{3N}$, where $C^{\prime}$ denotes the number of hand-aware conditioning channels. Note that partial layers of PVCNN are adapted for dual-stream conditioning and the details are discussed in the supplementary material.
During optimization, we use the MSE loss between model output $\epsilon_{\theta}(X_{t},t,z)$ and noise $\epsilon$:
$\mathcal{L}_{denoise}=\|\epsilon-\epsilon_{\theta}(X_{t},t,z)\|,\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (8)
Notice that the centroid prediction network is separately trained from $\mathcal{D}_{\theta}$. Thereby, the 3D object centroid $\mathcal{M}_{3d}$ in the hand wrist frame and the 2D object centroid $\mathcal{M}_{2d}$ in the normalized device coordinate (NDC) space are both supervised along with a 2D-3D projection loss. Overall, our total loss is defined as
$\begin{split}\mathcal{L}_{centroid}=&~{}\|\mathcal{M}_{3d}-\widehat{\mathcal{M}}_{3d}\|+\lambda_{1}\|\mathcal{M}_{2d}-\widehat{\mathcal{M}}_{2d}\|\\ &~{}+\lambda_{2}\|\mathcal{P}(\widehat{\mathcal{M}}_{3d})-\widehat{\mathcal{M}}_{2d}\|,\end{split}$ (9)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters, $\mathcal{P}$ is the transformation from the hand wrist frame to the NDC space and $\widehat{\bullet}$ refers to predicted results.
4 Experiments
Figure 3: Qualitative results on the ObMan [30] dataset. For each method and ground truth, we show the reconstruction results in the camera view (column 1) and a novel view (column 2).
4.1 Datasets and Setup
Datasets. We evaluate our method on three common benchmark datasets, including the synthetic ObMan [30] dataset and two real-world datasets, namely HO3D [29] and MOW [4]. ObMan [30] consists of 2,772 objects from 8 categories taken from ShapeNet [5], with 21K plausible grasps generated using GraspIt [46]. The dataset contains 141K frames for training and 6K frames for testing, which are rendered using Blender on top of random backgrounds. HO3D [29] includes 77,558 frames from 68 sequences, containing 10 different objects from the YCB dataset [3], which are manipulated by 10 different users. The hand and object pose annotations are obtained from multi-camera optimization procedures. MOW [4] consists of 442 images of 121 object templates, with in-the-wild source data collected from EPIC Kitchens [13, 14] and the 100 Days of Hands [62] datasets. We use the same training and testing split as iHOI [73]. Regarding hand pose prediction, we employ Hasson et al. [30] for the ObMan dataset and FrankMocap [60] for HO3D and MOW, in line with iHOI [73].
Evaluation Metrics. Following [73], we report the Chamfer Distance (CD) in mm and F-score at thresholds of 5mm (F-5) and 10mm (F-10) to compare with the state-of-the-art.
4.2 Evaluation on ObMan Dataset
We first present quantitative results on the large-scale synthetic ObMan dataset in Tab. 1. As can be seen, our approach surpasses all state-of-the-art methods by a large margin for F-score and Chamfer Distance. Specifically, compared with the current best-performing method gSDF [9], we achieve a relative improvement of 27.3 % and 13.6 % for the F-5 and F-10 metrics, respectively. Furthermore, we can reduce the CD by 73.5 % compared with iHOI and 57.8 % compared with HO, demonstrating that our reconstructed shape possesses significantly fewer outliers.
We also visualize the reconstructed objects along with the hands in Fig. 3. While SDF-based methods, including iHOI [73] and gSDF [9], tend to result in either over-smoothed and less detailed, or fragmented reconstructions, our approach is able to generate geometrically coherent point clouds with plausible details, even for thin objects and heavily occluded parts.
F-5 $\uparrow$ F-10 $\uparrow$ CD $\downarrow$
HO [30] 0.23 0.56 0.64
GF [37] 0.30 0.51 1.39
AlignSDF [8] 0.40 0.64 -
iHOI [73] 0.42 0.63 1.02
gSDF [9] 0.44 0.66 -
Ours 0.56 0.75 0.27
Table 1: Comparison with State of the Arts on ObMan [30]. F-score of 5mm and 10mm, Chamfer Distance (mm) metrics are utilized for evaluation.
HO3D MOW
Finetuning F-5 $\uparrow$ F-10 $\uparrow$ CD $\downarrow$ F-5 $\uparrow$ F-10 $\uparrow$ CD $\downarrow$
HO [30] 0.11 0.22 4.19 0.03 0.06 49.8
GF [37] 0.12 0.24 4.96 0.06 0.13 40.1
iHOI [73] 0.28 0.50 1.53 0.13 0.24 23.1
Ours 0.36 0.55 0.66 0.25 0.39 1.86
iHOI [73] 0.14 0.27 4.36 0.09* 0.17* 8.43*
Ours 0.22 0.35 1.34 0.22 0.33 3.15
Table 2: Results on real-world datasets HO3D [29] and MOW [4]. We report results for both finetuning (top) and zero-shot transfer (bottom).
Figure 4: Qualitative results on HO3D [29] (top) and MOW [4] (middle and bottom) datasets. For each method and ground truth, we show the reconstruction results in the camera view (column 1) and a novel view (column 2).
4.3 Evaluation on Real-World Datasets
Aside from the synthetic ObMan dataset, we also conduct experiments on the two real-world datasets HO3D [29] and MOW [4] to demonstrate our approach’s generalization capabilities for real-world scenarios. The model is respectively finetuned on the two datasets, starting from the ObMan pre-trained model as initialization. As shown in Tab. 2 (top), our approach achieves state-of-the-art results on HO3D as well as MOW. Concretely, regarding F-5 and F-10 metrics, we exhibit a noticeable average improvement of 19.3 % on the HO3D dataset and 77.4 % on the MOW dataset. Compared with the F-score metric, the Chamfer Distance metric is known to be more vulnerable to outliers [65, 73, 9]. Thus, the significant reduction of the CD metric illustrates that our approach is capable of producing significantly fewer outliers. The qualitative results in Fig. 4 also demonstrate the superiority of our approach.
Zero-shot transfer to HO3D and MOW. To further evaluate our zero-shot transfer abilities, we also directly apply our model, trained on the ObMan dataset, to HO3D and MOW without conducting any additional finetuning. The results in Tab. 2 (bottom) show that our method can again achieve a remarkable improvement on HO3D (57 % and 30 % in F-5 and F-10) and MOW (144 % and 94 % in F-5 and F-10), demonstrating the strength of HACD in synthetic-to-real generalization.
4.4 Ablation Studies
Row Method ObMan HO3D
F-5 $\uparrow$ F-10 $\uparrow$ CD $\downarrow$ F-5 $\uparrow$ F-10 $\uparrow$ CD $\downarrow$
A0 Ours 0.557 0.745 0.265 0.355 0.552 0.659
B0 A0 $\rightarrow$ w/o hand-aware conditioning 0.478 0.667 0.407 0.279 0.461 0.964
B1 A0 $\rightarrow$ w/o hand semantic embedding 0.510 0.694 0.373 0.325 0.496 0.809
B2 A0 $\rightarrow$ w/o hand articulation embedding 0.510 0.693 0.383 0.301 0.483 0.885
B3 A0 $\rightarrow$ w/ GCN hand embedding 0.523 0.713 0.297 0.338 0.537 0.522
B4 A0 $\rightarrow$ w/ global hand embedding 0.524 0.711 0.301 0.309 0.490 0.859
B5 A0 $\rightarrow$ w/o dual-stream conditioning 0.544 0.736 0.274 0.339 0.528 0.756
C0 B0 $\rightarrow$ w/o centroid fixing 0.439 0.609 0.650 0.268 0.452 0.997
C1 B0 $\rightarrow$ w/o centroid prediction 0.317 0.452 2.477 0.229 0.361 1.309
C2 C0 $\rightarrow$ Test with GT object centroid 0.450 0.667 0.362 0.289 0.469 0.932
C3 C0 $\rightarrow$ Test with GT object pose 0.499 0.703 0.343 0.307 0.485 0.841
D0 A0 $\rightarrow$ Predicted hand pose + noise $\sigma=0.1$ 0.556 0.745 0.265 0.348 0.536 0.690
D1 A0 $\rightarrow$ Predicted hand pose + noise $\sigma=0.5$ 0.517 0.707 0.319 0.321 0.504 0.837
E0 A0 $\rightarrow$ GT hand pose 0.596 0.768 0.236 0.372 0.568 0.601
E1 A0 $\rightarrow$ GT hand pose + noise $\sigma=0.1$ 0.572 0.761 0.262 0.358 0.556 0.642
E2 A0 $\rightarrow$ GT hand pose + noise $\sigma=0.5$ 0.535 0.726 0.325 0.314 0.523 0.691
F0 C2 $\rightarrow$ GT object centroid + noise $\sigma=0.1$ 0.442 0.647 0.341 0.281 0.459 1.002
F1 C2 $\rightarrow$ GT object centroid + noise $\sigma=0.5$ 0.411 0.601 0.495 0.236 0.422 1.228
Table 3: Ablation study. (a) Ablation on ObMan [30] and HO3D [29] datasets. (b) Robustness against hand occlusion. We analyze the patterns of the F-5, F-10, and CD metrics as a function of the object visibility ratio on the test set of ObMan [30].
Effectiveness of hand-aware conditioning. In Tab. 3 (a), we show the impact of our proposed hand-aware conditioning. Both the hand semantic embedding (B2 vs. B0, A0 vs. B1) and the geometric hand articulation embedding (B1 vs. B0, A0 vs. B2) result in improved performance. Noteworthy, we can surpass existing methods on the ObMan dataset even without using any hand-aware conditioning (B0).
We further noticed that the way of encoding the hand-aware conditional information has a significant impact on performance. In particular, we have explored two alternative strategies to encode the hand-object interaction. First, inspired by [20], we implement a GCN-based hand embedding, which utilizes a graph convolutional network [23] for feature extraction. We again apply the GCN hand embedding to the partially denoised point cloud at each step. Although the GCN hand embedding implicitly encodes the hand articulation, it does not encode the hand-object interaction. Therefore, simply applying such an embedding as conditioning does not improve performance (See B3 vs. A0). Our hand articulation embedding, on the other hand, applies the unique articulation-aware embedding to each point of the partially denoised point cloud, which explicitly encodes hand-object interaction, thus significantly benefiting the object reconstruction performance. Second, we also utilize standard PointNet [53] to encode 3D hand vertices into a global hand embedding, which is then served to the partially denoised point cloud. Again, without modeling the hand-object interaction, the global hand embedding falls short of providing sufficient semantic information about the hand-occluded object. Consequently, the predictions end up being inferior to our proposed hand semantic embedding (B4 vs. A0).
Effectiveness of dual-stream conditioning. Instead of integrating our unified semantic hand-object and geometric hand articulation embeddings by the proposed dual-stream architecture, we simply concatenate the embeddings. Note that we concatenate a zero matrix to the hand articulation embedding to meet the dimension of unified semantic hand-object embedding. The declined results (B5 vs. A0) demonstrate that dual-stream conditioning enhances reconstruction performance by independently processing semantic and geometric embeddings, ensuring their mutual independence.
Effectiveness of hand-constrained centroid fixing. In Tab. 3 (a), we illustrate the importance of the centroid fixing scheme. Our centroid fixing operation improves the precision and robustness of the semantic and geometric embedding, thus enhancing the performance (B0 vs. C0). Further, without the hand-constrained centroid prediction network, the diffusion model has to simultaneously learn the centroid deviation and object shape, which leads to clearly worse results (C1 vs. B0). Additionally, in C2 we report results when using the actual ground-truth object centroid and in C3 when using the ground-truth object pose. Without our centroid fixing scheme, the results remain non-comparable even when testing with the ground-truth object centroid (C2 vs. B0). This further demonstrates the effectiveness of our proposed centroid fixing scheme.
Robustness against hand pose prediction quality. To demonstrate that our method is able to deal with noisy hand pose estimates, we add various levels of Gaussian noise to predicted hand poses. Notice that adding $\sigma=0.1$ noise has almost no effect on the performance (D0 vs. A0) for both ObMan and HO3D. Further, despite a slight decrease in performance with $\sigma=0.5$ noise (D1 vs. A0), the results clearly show that our approach is robust towards weak hand pose predictions. Additionally, in E0 we explore the upper bound of our method using ground-truth hand poses, and in E1 / E2 we further demonstrate the robustness against noisy hand pose prediction.
Robustness against object centroid prediction quality. We also add different levels of Gaussian noise to the ground-truth object centroid (F0 / F1) to constitute the robustness of our method towards inaccurate object centroid.
Robustness against hand occlusion. To illustrate the robustness of our approach against hand occlusion, we split the test set of ObMan into groups according to the visible ratio of the object and compute the mean F-score and CD for each group. As shown in Tab. 3 (b), when the object is undergoing strong occlusion by the hand (object visible ratio $<$ 50%), iHOI suffers a significant decline in both F-5 and F-10 metrics, while the performance of our approach remains steady. The quantitative results of CD metric also demonstrate that aided by our proposed hand-aware conditioning and hand-constrained centroid fixing scheme, HACD is capable of modeling the uncertainties induced by hand occlusion.
5 Conclusion
In this paper, we present HACD, a novel hand-aware point cloud diffusion model for single-view hand-held object reconstruction. HACD does not require any object templates, category priors, or depth information, and exhibits superior performance in modeling the uncertainties induced by hand- and self-occlusion. In the core, we introduce hand-aware conditioning to model the semantic and geometric aspects of hand-object interactions. We further propose a hand-constrained centroid-fixing scheme, utilizing the estimated hand vertices to prevent the centroid of the partially denoised point cloud from diverging during diffusion and reverse processes. The experiments on ObMan, HO3D, and MOW datasets demonstrate that our approach surpasses existing methods in both synthetic and real-world scenarios.